# 2022-deformable-elastic-objects-UAVs

### v1

@@ Abstract
Aerial cooperative robotic manipulation of cable-suspended objects has been largely studied as it allows handling large and heavy objects, and cables offer multiple advantages, such as their low weight and cost efficiency. Multirotors have been typically considered, which, however, can be unsuitable for long-lasting manipulation tasks due to their scarce endurance. Hence, this work investigates whether non-stop flights are possible for maintaining constant the pose of cable-suspended objects. First, we show that one or two flying carriers alone cannot perform non-stop flights while maintaining a constant pose of the suspended object. Instead, we demonstrate that three flying carriers can achieve this task provided that the orientation of the load at the equilibrium is such that the components of the cable forces that balance the external force (typically gravity) do not belong to the plane of the cable anchoring points on the load. Numerical tests are presented in support of the analytical results.
@@ I Introduction
Aerial robotic object manipulation has been largely studied also thanks to its interesting applications, e.g., transportation and assembly, to name a few [1]. Among the several manipulator designs that have been proposed in the literature [2], lightweight manipulation tools such as simple low-cost cables have attracted a lot of attention [3], especially due to the flying vehicle payload limitations.
Uncrewed helicopters have been used in [4] for transportation of suspended objects, while multirotor UAVs (Uncrewed Aerial Vehicles) have been typically considered [5, 6, 7, 8, 9, 10]. In the literature, we find examples of a slung load suspended below a single multi-rotor [11, 12] or a small helicopter [13]. Two multirotors have been widely exploited, especially for the manipulation of bar-shaped objects [14, 15, 16]. Three is the minimum number of robots to control the full pose of a cable-suspended rigid body object [17]; cooperative manipulation of cable-suspended load below three multirotors have been studied, e.g., in [18, 5, 7, 19].
Compared to multirotor UAVs, fixed-wing UAVs have a longer flight endurance, as it has been shown, e.g., in [20]. However, fixed-wing UAVs cannot stop in mid-air, thus making seemingly impossible to keep the load in a static equilibrium while using such platforms as flying carriers. Furthermore these platforms are unable of Vertical Take Off and Landing (VTOL). With an intermediate flight endurance, convertible UAVs such as tail sitters have been proposed [20]. These are VTOL platforms that convert themselves to proceed flying horizontally as a fixed-wing platform. However, also these designs are obliged to follow non-stop trajectories to ensure an efficient flights.
@@
The theoretical discovery of such compatibility would pave the way to a completely new type of cable-suspended aerial systems which would enable a much more energy-efficient execution of aerial manipulation tasks in which long-distance transport is combined with fine manipulation of the object’s pose at the destination or in the presence of difficult intermediate passages, e.g., in cluttered environments. An abstract representation of the future scenario that would be possibly enabled by this study is shown in Figure 1.
@@
In this work, we take up the challenge posed by this intriguing problem and we study analytically whether there exist non-stop trajectories for the flying carriers that are compatible with the constraint of keeping the load at a constant static pose. We focus on the case in which the load must maintain a constant pose while the carriers move because it is obviously the most challenging case compared to the one in which also the load moves.
The work is organized as follows. First, in Sec. II, we introduce the notation and we state the problem formally. Hence, we answer the above-described research question in Sec. III, in which we analyze the cases of one, two, and three flying carriers. We show that 3 is the minimum number of flying carriers for which non-stop flights are possible and we derive the conditions for that to hold. In Sec. IV we show numerical examples in support of the theoretical results. Eventually, conclusions are drawn in Sec. V with an outline of future work.
@@ II-A Mathematical background
Consider ${\mathcal{F}_{W}=\{O_{W},\bm{x}_{W},\bm{y}_{W},\bm{z}_{W}\},}$ being a world-fixed frame where $O_{W}$ is the origin, and $\bm{x}_{W},\bm{y}_{W},\bm{z}_{W}$ are the x-, y-, and z- axes, respectively. Similarly, consider a frame attached to the Center of Mass (CoM) of a rigid body, which represents our manipulated object, and call it ${\mathcal{F}_{B}=\{O_{B},\bm{x}_{B},\bm{y}_{B},\bm{z}_{B}\}.}$ The position of $O_{B}$ in $\mathcal{F}_{W}$ is denoted with ${\bm{p}_{L}\in\mathbb{R}^{3}}$ and $\bm{R}_{L}\in SO(3)$ expresses the attitude of $\mathcal{F}_{B}$ w.r.t $\mathcal{F}_{W}$. The body angular velocity expressed in the body frame is indicated as ${}^{B}\bm{\omega}_{L}$.
The rigid body is manipulated through $n$ cables, each one of length $L_{i}>0$ and attached to point $B_{i}$ on the object, with ${}^{B}\bm{b}_{i}\in\mathbb{R}^{3}$ representing the constant position of $B_{i}$ expressed in $\mathcal{F}_{B}$ [^1] The left superscript expresses the reference frame, and if omitted it is considered equal to $W$ unless differently specified.. Let $\bm{q}_{i}\in S^{2}$ be the configuration of the $i^{\rm{th}}$ cable in $\mathcal{F}_{W}$, $T_{i}$ its tension, and
@@
The equations describing the dynamics of the transported object are
${m_{L}}\ddot{\bm{p}}_{L}$ $=-{m_{L}}g\bm{e}_{3}+\sum_{i=1}^{n}{\bm{f}_{i}}$ (2)
$\bm{J}_{L}{{}^{B}\dot{\bm{\omega}}}_{L}$ $=-\bm{J}_{L}{{}^{B}\bm{\omega}}_{L}\times{{}^{B}\bm{\omega}}_{L}+\sum_{i=1}^{n}{S(^{B}\bm{b}_{i})\bm{R}_{L}\bm{f}_{i}}$ (3)
@@
where ${m_{L}},\bm{J}_{L}$ are the mass and the rotational inertia of the rigid body load and $S(\star)$ indicates the skew operator that implements the cross product between two vectors.
@@
with $\star^{\dagger}$ indicating the Moore-Penrose pseudo-inverse, $\bm{N}\in\mathbb{R}^{3n\times m}$ being the nullspace projector of the grasp matrix $\bm{G}$ (with $m$ the dimension of the nullpsace of $\bm{G}$), and $\bm{\lambda}\in\mathbb{R}^{m}$ a vector of $m$ free parameters $\lambda_{1},\ldots,\lambda_{m}$ referred to as the internal forces [22].
For each load forced equilibrium or trajectory defined by the load state and the corresponding forcing input $\bm{W}$ we can obtain from (9) all the cable forces that are compatible with such load equilibrium/trajectory. These forces are parameterized by $\bm{\lambda}$. In turn, to each cable force corresponds a carrier position obtained by (5) substituting $\bm{q}_{i}=\bm{f}_{i}/T_{i}$. So we can conclude that for any given load equilibrium/trajectory there are $\infty^{m}$ compatible multi-UAV coordinated trajectories which are parameterized by the internal forces $\bm{\lambda}$ and are obtained by plugging (9) in (5).
@@ II-B Maintaining Load Static Equilibrium with non-stop flights
Departing from the previous literature, in this work we explore whether the internal forces can be exploited to allow the flying carriers to keep loitering with $\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}>0$ while the load pose is kept to a static pose.
Consider a static equilibrium of the suspended load characterized by position $\bar{\bm{p}}_{L}$ and orientation $\bar{\bm{R}}_{L}$. When the load is at a static equilibrium, the only way the cable forces can still change is thanks to a change of the internal forces $\bm{\lambda}$ since they are the only time-varying parameters in (9) in that case. For the purpose of analyzing the effect of changing internal forces we write them as the state of the following dynamical system:
@@
The first two equations in (10) impose static equilibrium of the load. The third equation, where $\underline{v}$ is a constant, is a lower bound to the norms of the flying carriers’ velocities to impose that they do not stop while maintaining the constantly desired pose of the object. Finally, the last expression in (10) imposes that each component of $\bm{\lambda}$ is bounded by certain upper and lower bounds, $\overline{\lambda},\underline{\lambda}$. This last constraint has clear practical reasons as internal forces of too large magnitude in (9) generate cable forces that will eventually break the system.
@@
By differentiating (1) w.r.t. time, the cable force variation (12) is also
@@
i.e., the force variation is composed of two orthogonal components: one is directed as $\bm{q}_{i}$ (along the cable), and one is $\dot{\bm{q}}_{i}$ (orthogonal to the cable).
@@ III Non-stop Flight Analysis
$\exists\underline{v}>0:\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}\;\forall t\Longrightarrow\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s\;\forall t$
@@
Proposition 1, when looked in reverse, provides a necessary condition for (10) to hold. Namely, in order to have all the carriers perform non-stop flights, all the norms of the cable forces derivatives must be lower bounded (i.e., if $\forall s>0,\ \|\dot{\bm{f}}_{i}(t)\|\ngeq s\Rightarrow\forall\underline{v}>0,\ \|\dot{\bm{p}}_{Ri}\|\ngeq\underline{v}$).
@@
$\bm{G}\in\mathbb{R}^{6\times 6}$ has a non-trivial nullspace of dimension $m=1$ and a basis of if is built from the vector along the direction through the points $B_{1}$ and $B_{2}$ [14]. We indicate the vector that constitutes such basis with $\bm{n}_{1}$, and $\lambda_{1}$ the corresponding scalar internal force. In particular, $\bm{N}={\bm{n}_{1}=\begin{bmatrix}\bm{b}_{1,2}^{\top}&\bm{b}_{2,1}^{\top}\end{bmatrix}^{\top}},$ where we indicate in general the unit vector in the direction $B_{i}B_{j}$ as ${\bm{b}_{i,j}=\frac{\bm{b}_{i}-\bm{b}_{j}}{|\bm{b}_{i}-\bm{b}_{j}|}}$.
@@
At the load equilibrium, $\bm{n}_{1}$ is constant by construction. Equation (12) leads to
@@
Assume that (10) holds. Then, for Proposition 1, $\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s$. For the considered case of 2 carriers, $\forall t>0,$ $\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s\Longleftrightarrow|\dot{\lambda}_{1}|\geq s.$ This follows from (17) considering that $\bm{b}_{i,j}$ has unit norm by construction. However, $|\dot{\lambda}_{1}|\geq s$ implies that $|\lambda_{1}|$ is a strictly increasing function of time and therefore is unbounded for $t\to\infty$. This contradicts the boundedness of $\lambda_{1}$ imposed in (10) and therefore (10) cannot hold as two of its conditions are in conflict. ∎
@@
In this section, we consider the case in which $n=3$ carriers are employed. Let us restrict the analysis to the non-degenerate cases in which the cable attachment points on the objects, $B_{i}$, are not all aligned to each other. In that degenerate case, the problem is similar to the 2-carrier scenario. In our study, we are going to assume that the minimum angle between two vectors $\bm{b}_{i,j}$ is $|\theta_{i,j}|\leq\bar{\theta}_{i,j}>0$.
For $n=3$, $\bm{\lambda}$ is a 3-dimensional vector of components $\lambda_{i}$, for $i\in\{1,2,3\}$, and a choice for $\bm{N}$ is expressed starting from the directions connecting the three points $B_{i}$ in pairs. In particular, it holds that [24, 22]:
@@
Equation (18) tells us that the components of each cable force generated by the internal forces are a couple of equal and opposite forces acting along the line connecting two cable attachment points.
Figure 2: Geometric representation of the cable force, its components, and its derivative. The grey vectors $\dot{\bm{f}}_{i}$ live in plane $P_{i}$, while its components $T_{i}\dot{\bm{q}}_{i},\dot{T}_{i}\bm{q}_{i}\notin P_{i}$ (in general). $\tilde{\bm{f}}_{i}$ is contained on the subspace $\tilde{P}_{i}$, which is centered at the end-point of $\bm{f}_{0i}$ and contains also the end-point of $\bm{f}_{i}$.
Figure 3: Representation of the plane including $\dot{\bm{f}}_{i}$ and its components along the directions parallel and perpendicular to the corresponding cable force. The minimum and maximum values of the angle between $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ are represented with dotted incident lines, and the circle of radius $s$ is the lower bound on the norm of $\dot{\bm{f}}_{i}.$ As a consequence, the norm of the component of $\dot{\bm{f}}_{i}$ on the direction orthogonal to the cable force, namely along $\dot{\bm{q}}_{i}$, is lower bounded by a positive quantity, $s^{\prime\prime}$.
Figure 4: Two carriers cannot maintain the pose of the object constant while performing non-stop flights. On the left, is a screenshot of an animation: the object is represented as a grey line, the cables are in black, and colored curves are the carrier’s paths. Carriers’ velocities and $\dot{\bm{\lambda}}$ are reported in the other two plots.
@@
For $n=3$, if the load’s attitude at the static equilibrium is such that $\forall i$ $\bm{f}_{0i}\notin{\rm{span}}\{\bm{b}_{i,j},\bm{b}_{i,k}\},$ with $i,j,k\in\{1,2,3\},$ $i\neq j\neq k$, then for $s,\underline{v}>0$, $\|{\dot{\bm{f}}_{i}\|\geq s\Rightarrow\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}.}$ $\forall i=1,\ldots,n$
@@
First of all, from (6) at the equilibrium, ${\|\dot{\bm{p}}_{Ri}\|=L_{i}\|\dot{\bm{q}}_{i}\|}$; hence, to prove the thesis of the proposition it is enough to show that $\|\dot{\bm{q}}_{i}\|\geq s^{\prime}$ with $s^{\prime}>0$. For the sake of a streamlined presentation we decided to avoid presenting all the mathematical details and to make use of explanatory figures that help the intuition of the reader being convinced about the validity of the proof.
@@
Please refer to to Fig. 2 to better understand the quantities defined in the following. Referring to (11) we know that ${\bm{f}_{i}=\bm{f}_{0i}+\tilde{\bm{f}}_{i}}$, where $\bm{f}_{0i}$ is obtained by the corresponding entries of $\bm{G}^{\dagger}\bm{W}$ and it is nonzero and constant in time, and $\tilde{\bm{f}}_{i}$ is obtained by the corresponding entries of $\bm{N}\bm{\lambda}$ and it is time-varying and belongs to the plane $P_{i}$ spanned by the vectors $\bm{b}_{i,j}$ and $\bm{b}_{i,k}$ with $j,k\neq i$ and $j\neq k$, according to (18). From the hypotheses of the proposition, we have that $\bm{f}_{0i}$ does not belong entirely to $P_{i}$ (apart from its endpoint). Since $\bm{f}_{0i}$ is constant and nonzero, also $\bm{f}_{i}$ does not belong entirely to $P_{i}$ (apart from its endpoint). On the other hand, from (12), we have that $\dot{\bm{f}}_{i}$ belongs entirely to $P_{i}$. As a consequence, thanks also to the hypothesis that $\dot{\bm{f}}_{i}\neq\bm{0}$, we obtain that $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ are never parallel to each other (two parallel vectors must belong to the same plane). In other words, the angle between each cable force and its derivative is $0<\angle\bm{f}_{i},\dot{\bm{f}}_{i}<\pi.$
@@
Because of the boundedness of each component of $\bm{\lambda}$, the endpoint of $\bm{f}_{i}$ is constrained inside a bounded set denoted with $\tilde{P}_{i}\in P_{i}$, the dark grey region in Figure 2. $\tilde{P}_{i}$ is centered around the end-point of $\bm{f}_{0i}$ and contains the vector $\tilde{\bm{f}}_{i}.$ For all the $\bm{f}_{i}$, we have a minimum and maximum of the angle $\angle\bm{f}_{i},\dot{\bm{f}}_{i}$ when $\dot{\bm{f}}_{i}$ varies in $\tilde{P}$, which we denote as $\alpha(\bm{f}_{i})$ and $\beta(\bm{f}_{i})$ respectively. Furthermore, we denote with $\underline{\alpha}$ and $\overline{\beta}$ the minimum and maximum of of $\alpha(\bm{f}_{i})$ and $\beta(\bm{f}_{i})$ when the end point of $\bm{f}_{i}$ varies in $\tilde{P}_{i}$, respectively. Since those functions are continuous and the set $\tilde{P}_{i}$ is bounded such minimum and maximum exist and therefore, for what has been previously said, they are such that $0<\underline{\alpha}\leq\overline{\beta}<\pi.$

### v3

@@ Abstract
This work answers positively the question whether non-stop flights are possible for maintaining constant the pose of cable-suspended objects. Such a counterintuitive answer paves the way for a paradigm shift where energetically efficient fixed-wing flying carriers can replace the inefficient multirotor carriers that have been used so far in precise cooperative cable-suspended aerial manipulation. First, we show that one or two flying carriers alone cannot perform non-stop flights while maintaining a constant pose of the suspended object. Instead, we prove that three flying carriers can achieve this task provided that the orientation of the load at the equilibrium is such that the components of the cable forces that balance the external force (typically gravity) do not belong to the plane of the cable anchoring points on the load. Numerical tests are presented in support of the analytical results.
@@ I Introduction
Aerial robotic object manipulation has been largely studied also thanks to its interesting applications, e.g., transportation and assembly, to name a few [1]. Among the several manipulator designs proposed in the literature [2], simple lightweight and low-cost cables have attracted a lot of attention [3], especially due to the flying vehicle payload limitations.
Uncrewed helicopters have been used in [4] for transportation of suspended objects, while multirotor UAVs (Uncrewed Aerial Vehicles) have been typically considered [5, 6, 7, 8, 9, 10]. In the literature, we find examples of a slung load suspended below a single multi-rotor [11, 12] or a small helicopter [13]. Two multirotors have been widely exploited, especially for the manipulation of bar-shaped objects [14, 15, 16]. Three is the minimum number of robots to control the full pose of a cable-suspended rigid body object [17], and three multirotors have been studied, e.g., in [18, 5, 7, 19].
Compared to multirotor UAVs, fixed-wing UAVs have a longer flight endurance, as it has been shown, e.g., in [20]. However, fixed-wing UAVs cannot stop in mid-air, thus making seemingly impossible to keep the load in a static equilibrium while using such platforms as flying carriers. Furthermore, these platforms are unable to Vertical Take Off and Landing (VTOL). With an intermediate flight endurance, convertible UAVs such as tail sitters have been proposed [20]. These are VTOL platforms that convert themselves to proceed flying as a fixed-wing platform. However, also these designs are obliged to follow non-stop trajectories to ensure an efficient flight.
@@
The theoretical discovery of such compatibility would pave the way to a completely new type of cable-suspended aerial systems, which would enable a much more energy-efficient execution of aerial manipulation tasks: long-distance transport would be combined with fine manipulation of the object’s pose at the destination or in the presence of difficult intermediate passages, e.g., in cluttered environments. Figure 1 shows an abstract representation of the possible future scenario that would be enabled by this study.
@@
In this work, we consider the challenges posed by the described intriguing problem and we study analytically whether there exist non-stop trajectories for the flying carriers that are compatible with the constraint of keeping the load at a constant static pose. We focus on the case in which the load must maintain a constant pose because it is obviously the most challenging case compared to the one in which also the load moves. The main contributions of the work are as follows: (i) we formalize the problem in mathematical terms; (ii) we prove that with one and two carriers, non-stop flights are not achievable in the considered manipulation scenario; (iii) we prove that three is the minimum number of flying carriers for which non-stop flights are possible and (iv) we derive the conditions for that to hold, highlighting degenerate cases in which non-stop flights are prevented; (v) we support the theoretical results through numerical simulation.
The work is organized as follows. First, in Sec. II, we introduce the notation and the problem. Hence, we answer the above-described research question in Sec. III. In Sec. IV, we show numerical examples. Eventually, conclusions are drawn in Sec. V with an outline of future work.
@@ II-A Mathematical background
Consider ${\mathcal{F}_{W}=\{O_{W},\bm{x}_{W},\bm{y}_{W},\bm{z}_{W}\},}$ being a world-fixed frame where $O_{W}$ is the origin, and $\bm{x}_{W},\bm{y}_{W},\bm{z}_{W}$ are the x-, y-, and z- axes, respectively. Similarly, consider a frame attached to the Center of Mass (CoM) of a rigid body, which represents the manipulated object, and denote it with ${\mathcal{F}_{B}=\{O_{B},\bm{x}_{B},\bm{y}_{B},\bm{z}_{B}\}.}$ The position of $O_{B}$ in $\mathcal{F}_{W}$ is denoted with ${\bm{p}_{L}\in\mathbb{R}^{3}}$ and $\bm{R}_{L}\in SO(3)$ expresses the attitude of $\mathcal{F}_{B}$ w.r.t. $\mathcal{F}_{W}$. The body angular velocity expressed in the body frame is indicated as ${}^{B}\bm{\omega}_{L}$.
The rigid body is manipulated through $n$ cables, each one of length $L_{i}>0$ (with $i=1\ldots,n$) and attached to point $B_{i}$ on the object, with ${}^{B}\bm{b}_{i}\in\mathbb{R}^{3}$ representing the constant position of $B_{i}$ expressed in $\mathcal{F}_{B}$ [^1] The left superscript, $W$ if omitted, expresses the reference frame.. Let $\bm{q}_{i}\in S^{2}$ be the configuration of the $i^{\rm{th}}$ cable in $\mathcal{F}_{W}$, $T_{i}$ its tension, and
@@
The dynamics of the transported object is described by
${m_{L}}\ddot{\bm{p}}_{L}$ $=-{m_{L}}g\bm{e}_{3}+\textstyle\sum_{i=1}^{n}{\bm{f}_{i}}$ (2)
$\bm{J}_{L}{{}^{B}\dot{\bm{\omega}}}_{L}$ $=-\bm{J}_{L}{{}^{B}\bm{\omega}}_{L}\times{{}^{B}\bm{\omega}}_{L}+\textstyle\sum_{i=1}^{n}{S(^{B}\bm{b}_{i})\bm{R}_{L}\bm{f}_{i}}$ (3)
@@
where ${m_{L}},\bm{J}_{L}$ are the mass and the rotational inertia of the rigid body load and $S(\star)$ indicates the skew-symmetric operator implementing the cross product between two vectors.
@@
being $\star^{\dagger}$ indicating the Moore-Penrose pseudo-inverse, $\bm{N}\in\mathbb{R}^{3n\times m}$ being the nullspace projector of the grasp matrix $\bm{G}$ (with $m$ the dimension of the nullpsace of $\bm{G}$), and $\bm{\lambda}\in\mathbb{R}^{m}$ a column vector containing $m$ free parameters $\lambda_{1},\ldots,\lambda_{m}$ referred to as the internal forces [22].
For each load forced equilibrium or trajectory defined by the load state and the corresponding forcing input $\bm{W}$ we can obtain from (9) all the cable forces that are compatible with such load equilibrium/trajectory. These forces are parameterized by $\bm{\lambda}$. In turn, to each cable force corresponds a carrier position obtained by (5) substituting $\bm{q}_{i}=\bm{f}_{i}/T_{i}$, i.e., $\bm{q}_{i}=\bm{f}_{i}/\|\bm{f}_{i}\|$. So we can conclude that for any given load equilibrium/trajectory there are $\infty^{m}$ compatible multi-UAV coordinated trajectories which are parameterized by the internal forces $\bm{\lambda}$ and are obtained by plugging (9) in (5).
@@ II-B Maintaining Load Static Equilibrium with non-stop flights
Departing from the previous literature, in this work we explore whether the internal forces can be exploited to allow the flying carriers to keep loitering with $\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}>0$ while the load is kept to a static pose i.e., $\dot{\bm{p}}_{L}={{}^{B}\bm{\omega}}_{L}=\bm{0}$.
Consider a static equilibrium of the suspended load characterized by position $\bar{\bm{p}}_{L}$ and orientation $\bar{\bm{R}}_{L}$. When the load is at static equilibrium, the only way the cable forces can still change is thanks to a change of the internal forces $\bm{\lambda}$ since they are the only time-varying parameters in (9) in that case. For analyzing the effect of changing internal forces we write them as the state of the following dynamical system:
@@
The first two equations in (10) impose static equilibrium of the load. The third equation, where $\underline{v}$ is a constant, is a lower bound to the norms of the flying carriers’ velocities to impose that they do not stop while maintaining the desired constant pose of the object. Finally, the last expression in (10) imposes that each component of $\bm{\lambda}$ is bounded by certain upper and lower bounds, $\overline{\lambda},\underline{\lambda}$. This last constraint has clear practical reasons as internal forces of too large magnitude in (9) generate cable forces that will eventually break the system.
@@
By differentiating (1) w.r.t. time, the cable force variation (12) it holds
@@
i.e., the force variation is composed of two orthogonal components: one is directed as $\bm{q}_{i}$ (along the cable), and one as $\dot{\bm{q}}_{i}$ (orthogonal to the cable).
@@ III Non-stop Flight Analysis
$\exists\underline{v}>0\text{~{}s.t.~{}}\forall t\;\|\dot{\bm{p}}_{Ri}(t)\|\geq\underline{v}\Longrightarrow\exists s>0\text{~{}s.t.~{}}\forall t\;\|\dot{\bm{f}}_{i}(t)\|\geq s$
@@
Proposition 1 tells us that the lower boundedness of the norms of the cable forces derivatives is a necessary condition for (10) to hold. Namely, in reverse, if the norms of the cable forces derivatives are not lower bounded, then non-stop flights are impossible (i.e., $\nexists s>0\text{~{}s.t.~{}}\forall t\ \|\dot{\bm{f}}_{i}(t)\|\geq s\Rightarrow\nexists\underline{v}>0\text{~{}s.t.~{}}\forall t\ \|\dot{\bm{p}}_{Ri}\|\geq\underline{v}$).
@@
$\bm{G}\in\mathbb{R}^{6\times 6}$ has a non-trivial nullspace of dimension $m=1$ and a basis of it is built from the vector along the direction connecting the cable attachment points on the load, $B_{1}$ and $B_{2}$ [14].
In particular, $\bm{N}={\begin{bmatrix}\bm{b}_{1,2}^{\top}&\bm{b}_{2,1}^{\top}\end{bmatrix}^{\top}},$ where we indicate in general the unit vector in the direction $B_{i}B_{j}$ as ${\bm{b}_{i,j}=\frac{\bm{b}_{i}-\bm{b}_{j}}{|\bm{b}_{i}-\bm{b}_{j}|}}$, with $(i,j)\in\{(1,2),(2,1)\}$.
@@
We proceed by contradiction. At the load equilibrium, $\bm{b}_{i,j}$ is constant by construction. Equation (12) leads to
@@
Assume that (10) holds. Then, for Proposition 1, $\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s$. For the considered case of 2 carriers, $\forall t>0,$ $\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s\Longleftrightarrow|\dot{\lambda}_{1}|\geq s.$ This follows from (17) considering that $\bm{b}_{i,j}$ has unit norm by construction. Considering that $|\dot{\lambda}_{1}|\geq s$ must hold $\forall t$ and that $\dot{\lambda}_{1}$ is a continuous function of $t$, then either $\dot{\lambda}_{1}\geq s$ $\forall t$ and thus $\lambda_{1}$ is strictly increasing, or $\dot{\lambda}_{1}\leq-s$ $\forall t$ and thus $\lambda_{1}$ is strictly decreasing. In both cases, $\lambda_{1}$ is unbounded for $t\to\infty$. This contradicts the boundedness of $\lambda_{1}$ imposed in (10) and therefore (10) cannot hold as two of its conditions are in conflict. ∎
@@
In this section, we consider the case in which $n=3$ carriers are employed. Let us restrict the analysis to the non-degenerate cases in which the cable attachment points on the objects, the $B_{i}$’s, are not all aligned with each other. In that degenerate case, the problem is similar to the 2-carrier scenario. In our study, we are going to assume that any two vectors $\bm{b}_{i,j}$ and $\bm{b}_{i,k}$ with $(i,j,k)\in\{(1,2,3),(2,3,1),(3,1,2)\}$ are sufficiently misaligned. In particular, there exists $\theta_{\text{min}}>0$ that lower bounds any possible angle generated by the lines that the pairs $\bm{b}_{i,j}$, $\bm{b}_{i,k}$ belong to.
For $n=3$, $\bm{\lambda}=[\lambda_{1}\,\lambda_{2}\,\lambda_{3}]^{\top}$, and a choice for $\bm{N}$ is expressed starting from the directions connecting the three points $B_{i}$ in pairs. In particular, it holds that [24, 22]:
@@
Equation (18) tells us that the components of each cable force generated by the internal forces are a couple of equal and opposite forces acting along the line connecting two cable attachment points. In the following, we show that the necessary condition of Proposition 1 becomes sufficient for $n=3$ under certain assumptions.
Figure 2: Geometric representation of the cable force, its components, and its derivative. The grey vector $\dot{\bm{f}}_{i}$ belongs to the plane $P_{i}$, while its components $T_{i}\dot{\bm{q}}_{i},\dot{T}_{i}\bm{q}_{i}\notin P_{i}$ (in general). $\tilde{\bm{f}}_{i}$ is contained in the subset $\tilde{P}_{i}$.
Figure 3: Representation of the plane including $\dot{\bm{f}}_{i}$ and its components along the directions parallel and perpendicular to the corresponding cable force. The minimum and maximum values of the angle between $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ are represented with dotted half-lines, incident in the middle point, and the circle of radius $s$ is the lower bound on the norm of $\dot{\bm{f}}_{i}$. As a consequence, the norm of the component of $\dot{\bm{f}}_{i}$ on the direction orthogonal to the cable force, namely along $\dot{\bm{q}}_{i}$, is lower bounded by a positive quantity $\underline{v}^{\prime\prime}$.
Figure 4: On the left: two carriers (black dots) cannot maintain the pose of the object (in grey) while performing non-stop flights; the cables are in black, and colored curves are the carrier’s paths. Carriers’ velocities and $\dot{\bm{\lambda}}$ are reported in the other two plots.
@@
For $n=3$, if the load’s attitude at the static equilibrium is such that $\forall i$ $\bm{f}_{0i}\notin{\rm{span}}\{\bm{b}_{i,j},\bm{b}_{i,k}\}$, with $i,j,k\in\{1,2,3\},$ $i\neq j\neq k\neq i$ then the following is true ${\forall i=1,2,3}$
${\exists s>0\text{ s.t }}\forall t\ \|{\dot{\bm{f}}_{i}(t)\|\geq s\;\Rightarrow\;\exists\underline{v}>0\text{ s.t }\forall t\ \|\dot{\bm{p}}_{Ri}(t)\|\geq\underline{v}}$
@@
First of all, from (6) at the equilibrium, ${\|\dot{\bm{p}}_{Ri}\|=L_{i}\|\dot{\bm{q}}_{i}\|}$; hence, to prove the thesis of the proposition it is equivalent to show that $\|\dot{\bm{q}}_{i}\|\geq\underline{v}^{\prime}$ for a certain $\underline{v}^{\prime}>0$. For the sake of a streamlined presentation, we avoid presenting all the mathematical details and to make use of explanatory figures that help the intuition of the reader being convinced about the validity of the proof.
@@
Please refer to Fig. 2 to better understand the quantities defined in the following. Referring to (11) we know that ${\bm{f}_{i}=\bm{f}_{0i}+\tilde{\bm{f}}_{i}}$, where $\bm{f}_{0i}$ is obtained by the corresponding entries of $\bm{G}^{\dagger}\bm{W}$ and it is nonzero and constant in time, and $\tilde{\bm{f}}_{i}$ is obtained by the corresponding entries of $\bm{N}\bm{\lambda}$ and it is time-varying and belongs to the plane $P_{i}$ spanned by the vectors $\bm{b}_{i,j}$ and $\bm{b}_{i,k}$ with $j,k\neq i$ and $j\neq k$, according to (18). From the hypotheses of the proposition, we have that $\bm{f}_{0i}$ does not belong to $P_{i}$. Since $\bm{f}_{0i}$ is constant and nonzero, also $\bm{f}_{i}$ does not belong to $P_{i}$. On the other hand, from (12), we have that $\dot{\bm{f}}_{i}$ belongs to $P_{i}$. As a consequence, thanks also to the hypothesis that $\dot{\bm{f}}_{i}\neq\bm{0}$, we obtain that $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ are never parallel to each other (two parallel vectors must belong to the same plane). In other words, the angle between each cable force and its derivative is $0<\angle(\bm{f}_{i},\dot{\bm{f}}_{i})<\pi.$
@@
Because of the boundedness of each component of $\bm{\lambda}$, the endpoint of $\bm{f}_{i}$ is constrained inside a bounded set denoted with $\tilde{P}_{i}\subset P_{i}$, the dark grey region in Figure 2. $\tilde{P}_{i}$ is centered around the end-point of $\bm{f}_{0i}$ and contains the vector $\tilde{\bm{f}}_{i}.$ For each $\bm{f}_{i}$, with $i=1,2,3$, we have a minimum and maximum of the angle $\angle(\bm{f}_{i},\dot{\bm{f}}_{i})$ when $\dot{\bm{f}}_{i}$ varies in $\tilde{P}$, which we denote as $\alpha(\bm{f}_{i})$ and $\beta(\bm{f}_{i})$ respectively. We denote with $\underline{\alpha}$ and $\overline{\beta}$ the minimum and maximum of $\alpha(\bm{f}_{i})$ and $\beta(\bm{f}_{i})$ when the end point of $\bm{f}_{i}$ varies in $\tilde{P}_{i}$, and $i$ varies in $\{1,2,3\}$, respectively. Since those functions are continuous and the set $\tilde{P}_{i}$ is bounded such minimum and maximum exist and therefore, for what has been previously said, they are such that $0<\underline{\alpha}\leq\overline{\beta}<\pi.$